OpenStudy (anonymous):

Abstract Algebra Question: Let R be an equivalence relation on the set of real differentiable functions defined by fRg iff f and g has the same first derivative, i.e. f' = g'. Determine the equivalence classes of f for each f element of R.

OpenStudy (kinggeorge):

The equivalence class of f would just be the set of functions such that every function in the set is equal to f plus a constant term (The derivative of a constant term is 0).

OpenStudy (anonymous):

so should i define f. that means should i specify f in order to find the classes?

OpenStudy (kinggeorge):

f is just a real differentiable function. The most important thing is to define what a constant is. In this case, you should define the constant by saying it's a continuous differentiable function such that it's first derivative is 0.

OpenStudy (anonymous):

so can i say x^2/R= X^2+C, C= is a constant

OpenStudy (anonymous):

or ax^2+bx+C , a,b,c are constant.. can i define the classes like that?

OpenStudy (kinggeorge):

Partially. The only problem with defining classes like that, is you would have an infinite number of definitions, and you wouldn't include function like sin(x) and \(e^x\). Rather, I might recommend a definition similar to the following...

OpenStudy (kinggeorge):

Let \(f, g, h, a\) be continuous differentiable functions on the set of real numbers, and define an equivalence relation R such that \(f\text{R} g \;\Leftrightarrow \; f'=g'\). Then \(h\) is in the equivalence class of \(f\) iff \(h=f+a\) where \(a\;'=0\)

OpenStudy (anonymous):

mmm.. sounds good.. i think ill never come up with that.. :( why?

OpenStudy (kinggeorge):

It just takes practice. When I first started doing this, it took me a while to get used to it. Just make sure to proofread your definitions.

OpenStudy (anonymous):

thank you very much

OpenStudy (kinggeorge):

You're welcome.